OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 36°?

OpenStudy (anonymous):

360/36=n (The sum of exterior angles is 360°/sum of one exterior angle) Total interior angle= (n-2)*180

OpenStudy (anonymous):

oh yes! thank you

OpenStudy (anonymous):

you're welcome! :)

OpenStudy (anonymous):

Welcome to OpenStudy!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

!

OpenStudy (anonymous):

:)